OpenStudy (anonymous):

Divide using Long Division: (x^3 + x^2 +2x -4) divided by (x-1) A. x^2 + 2x + 4 B. x^2 + x - 4 C. x^2 + 4

OpenStudy (anonymous):

it is very hard to write long division here do you need the method or the answer?

OpenStudy (anonymous):

um watever one u wnt to give me..i'll take both or if u wnt u can just give me the answer

OpenStudy (anonymous):

since you have the possible answers it would be easiest to multitply and see which ones work

OpenStudy (anonymous):

um i dnt get it.

OpenStudy (anonymous):

it is \[x^2+2x+4\]

OpenStudy (anonymous):

x^2 x - 1 | x^3 + x^2 +2x -4 x^3 - x^2 ________ 2x^2 + 2x etc

OpenStudy (anonymous):

or \[x^3 + x^2 +2x -4)= (x-1) \times (\text{something})\] and you can figure out what the something has to be

OpenStudy (anonymous):

okay yea i get it now thnk u satellite lol